1/4x=1/20x+1 how do you solve this step by step
Answer by stanbon(75887)

(1/4)x = (1/20)x + 1
Multiply thru by 20 to get:
5x = x + 20
4x = 20
x = 5
------------
Cheers,
Stan H.
============